OpenStudy (anonymous):

Find an equation of the circle that satisfies the stated conditions. Endpoints of a diameter A(2, −5) and B(−6, 3)

OpenStudy (anonymous):

Find the midpoint of the diameter, which is -2, -1 and there you will have the center of your circle. From there, calculate the radius from finding the distance between the center of the circle -2,-1 and the endpoint of a diameter. Distance from (-2,-1) and (-6,3) = 4sqrt(2), which is your radius The equation of a circle is (x-h)^2 + (y-h)^2 = r^2 Now plug in what we know (x+2)^2 + (y + 1) ^2 = 32

OpenStudy (anonymous):

(x-h)^2 + (y - k)^2 = r^2 is the equation for a circle by the way

OpenStudy (anonymous):

i know that i just didnt know how to plug the numbers in

OpenStudy (anonymous):

plug the numbers in where?